OpenStudy (anonymous):

Which recursive formula can be used to generate the sequence shown, where f(1) = 5 and n > 1? 5,–1, –7, –13, –19, ...

OpenStudy (jango_in_dtown):

see difference of each term here is -6 from the previous one. so its an arithmetic progression sequence given by f(n+1)=f(n)-6,n>1 and f(1)=5

OpenStudy (jango_in_dtown):

@tristan1715

OpenStudy (anonymous):

thak you @jango_IN_DTOWN but thats not an anwser choice lol..

OpenStudy (anonymous):

ohhh jk THANK YOUUU @jango_IN_DTOWN

OpenStudy (jango_in_dtown):

give the choices.. but i am pretty sure the problem is solved